# Examining dark neutron decay using f-mode gravitational waves

### v1

@@ Abstract
Furthermore, fundamental mode (f-mode) oscillations have been analyzed, which serve as a diagnostic tool for probing neutron star interiors. Both first- and second-order f-mode frequencies follow universal relations, reinforcing their applicability for constraining dense matter properties. It has been shown that neutron stars composed of nucleons-only and self-interacting dark matter exhibit a universal behavior in damping time and angular frequency, whereas strange matter and non-self-interacting dark matter deviate from this trend.
These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer. This study provides a pathway for testing dark matter models through gravitational wave asteroseismology, bridging astrophysical observations and fundamental particle physics.
@@ 1 Introduction
The hypothesis of self-interacting dark matter (SIDM) has gained prominence due to its potential to address discrepancies within the standard collisionless cold dark matter (CCDM) paradigm, alongside its theoretical plausibility within extensions of the Standard Model. Specifically, the CCDM model faces challenges such as the observed flat density profiles in dwarf galaxy cores, contrasting with the cuspy profiles predicted by simulations [1, 2], the overabundance of simulated satellite galaxies compared to observed counts [1, 3, 4], and the ”too big to fail” problem, wherein simulations predict massive, star-forming dwarf galaxies that remain undetected [5].
@@
Additionally, the neutron lifetime anomaly, where beam experiments yield longer lifetimes than bottle experiments [11, 12, 13, 14], suggests a potential decay channel into undetected dark sector particles [15, 16, 17, 18]. This has motivated the exploration of neutron decay into dark matter as a potential resolution. The neutron decay scenario, proposed by Fornal and Grinstein [17], posits that neutrons decay into a dark fermion (), potentially explaining the observed neutron lifetime discrepancy. This hypothesis implies that neutron stars, composed primarily of neutrons, could serve as natural laboratories for studying dark matter interactions. If neutrons decay into dark matter within these stars, the accumulated dark matter could significantly influence their properties.
@@
In this study, well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear EoS, have been revisited [27, 28]. This investigation has been extented to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
In essence, f-mode oscillations provide a unique window into the microphysics of compact stars. For dark matter candidates, this approach opens a novel and testable pathway to constrain their properties via gravitational wave asteroseismology.
This paper has been divided into different sections. The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints. Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena. Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications. Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity. Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations. Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships. This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms. Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance. Finally, Section 8 will conclude by focusing on the f- mode oscillation frequencies of neutron stars and their importance for probing their internal structure and for gravitational wave astronomy.
@@ 2 Neutron decay into dark matter
This hypothesis posits that weakly interacting dark matter particles remain undetected in beam experiments but contribute to the total decay rate in bottle experiments. This alternative to neutron-mirror neutron oscillation [15, 16] necessitates a dark fermion with a mass (mχ) nearly degenerate with the neutron
Within the range [15, 16]. The accompanying boson cannot be a photon [38, 39]. Furthermore, this dark decay hypothesis has been suggested as a potential resolution to the reactor anti-neutrino anomaly [39]. Notably, it imposes constraints on neutron star properties. If these dark fermions are non-interacting, the maximum neutron star mass is predicted to be below 0.7 M⊙ [40, 41, 42, 43, 44], contradicting observed neutron star masses exceeding 2 M⊙ [45, 46, 47, 48]. Therefore, a repulsive self-interaction among dark fermions is required to reconcile this discrepancy.
This study investigates the implications of this dark decay channel on neutron star properties. This study takes forward the previous findings and constrains the properties of the dark matter using the gravitational wave, which can be very crucial in determining the interior of the neutron star.
@@ 3 Neutron Stars
The hypothesis that neutrons may decay into DM suggests that a substantial number of dark matter particles may accumulate within neutron stars, potentially leading to observable alterations in their macroscopic properties [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 43, 94, 95, 96, 97, 98, 99, 100, 101]. To model the nuclear matter inside neutron stars with precision, the equation of state (EoS) based on the quark-meson coupling (QMC) model has been used [102, 103, 104, 105]. This model is rooted in a relativistic mean field. This framework, grounded in relativistic mean-field theory, effectively represents nucleon interactions through meson exchange. In this study, I deliberately limit the hadronic composition to nucleons alone, omitting hyperons and strange quark matter. This simplification allows us to focus specifically on the effects of neutron-to-dark matter decay within a well-defined nuclear setting, facilitating a more precise analysis of its influence on neutron star properties. Through this approach, I systematically examine how dark matter accumulation affects observable features of neutron stars, offering valuable insights into dark matter’s fundamental characteristics.
@@ 4 Equation of State
To determine the equation of state, the QMC model has been adapted, which has shown promising results in [82, 106, 107, 108]. The quark-meson coupling (QMC) model was first introduced by Guichon and later expanded upon by Guichon, Thomas, and their collaborators. In this framework, nucleons are modeled as clusters of three quarks confined within an MIT bag, emphasizing their internal structure—unlike other models that treat nucleons as point-like particles [109]. The interactions between baryons arise from meson exchange, with the mesons coupling dynamically to the confined quarks. A key feature of this model is the strong scalar mean field, which induces significant modifications in the properties of bound baryons. As proposed by Fornal and Grinstein [17, 16], if neutron stars decay into dark matter. Dark matter must be taken into account when formulating an equation of state. At ()-equilibrium, the chemical equilibrium equations governing the composition of matter are:
@@
where is the mass of corresponding particle (subscripted). The nucleon effective mass, , given by
@@
The equation of state (EoS) describes the relationship between pressure () and energy density (). A generic form including dark neutron contributions can be written as:
where and represent the baryonic and dark neutron contributions, respectively. A common parameterization for dark matter influence is given by:
where is the dark matter equation of state parameter. The nucleons-only EoS is taken from [105] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
where the energy density of the strange matter and is the bag constant, taken to be 1014gm/fm3.
@@ 6 Relationship between energy density and pressure
Figure 1 illustrates the equation of state of different models selected in this study. For the nucleon-only equation of state, the QMC model is selected. The equation of state describing neutron decay into dark matter is presented for different DM-DM self-interactions. Various equations of state are considered, each corresponding to a different DM-DM self-interaction. For comparison, strange matter EoS is also given. Fig. 1 shows that the nucleon-only equation of state is the stiffest among the selected models, whereas increasing DM-DM self-repulsion makes the equation of state stiffer as the self-interaction increases the equation of state neutron decay into DM comes closer to the nucleon-only EoS. Here the self-interaction of dark matter is thought to resemble the neutron- coupling, represented by (). It is defined as:
where is the coupling constant, and is the mass of the interchange particle. Various values of are selected in this study, where the number next to represents the strength of the DM-DM particle interaction. A higher value indicates stronger DM-DM self-interaction, influencing the overall dynamics and structural properties of the system.
The strange matter EoS poorly represents the properties of nuclear matter at lower energy densities. Therefore, below 500 MeV/fm3, the strange matter EoS matches with the nucleon-only EoS [105, 107].
@@
Fig. 2 represents the mass and radius of neutron stars based on different EoSs. Ref. [115, 116, 117] report the observation of neutron stars of a mass of exceeding 2 M⊙, indicating that any viable model must support a maximum neutron star mass above this threshold 2 M⊙. The nucleon-only EoS produces a neutron star of maximum mass greater than 2 M⊙, the strange matter EoS produces a neutron star of maximum mass close to 2 M⊙, while different EoSs based on different DM-DM self-interactions produce a neutron star of less than 2 M⊙. The non-self-interactive ( = 0 fm2) DM EoS produces a neutron star of mass close to 0.7 M⊙ which is well below the acceptable value. When the DM-DM self-interaction . The EoS produces a neutron star of 2 M⊙. Therefore, acceptable self-interaction strength is 26 fm2.
@@ 7 Universalities
The universalities refer to empirical relations that link the fundamental oscillation modes of compact objects, such as neutron stars and dark stars, to their macroscopic properties like mass, radius, and compactness. These universal relations are valuable because they provide a model-independent way to infer stellar characteristics from gravitational wave observations [118, 113]. Traditional neutron stars follow well-established scaling laws for f-mode frequencies and damping times, but recent studies show that dark matter admixed compact stars deviate from these relations, requiring new frameworks to describe their oscillations [119, 120]. By leveraging f-mode gravitational wave signals, astrophysicists aim to refine equations of state and probe exotic forms of matter beyond the Standard Model.
@@ 7.1 Damping time
Reference [27] introduced a first-order universal relation describing the damping time () of f-mode oscillations. Their estimate was based on the quadrupole radiation formula, and the corresponding result is illustrated in Figure 3.
This equation expresses the characteristic damping time () of an oscillating astrophysical object, showing its relationship to the oscillation energy and the power emitted through gravitational waves (GW). The notation suggests that the damping time is proportional to the total oscillation energy divided by the power emitted as gravitational waves. The duration of oscillations in an astrophysical system depends on how efficiently it emits gravitational waves, with larger stars generally retaining their oscillations longer. Building on this, Lioutas and Stergioulas [121] later refined the relation by including higher-order corrections. The results illustrated in Fig. (3) demonstrate the first and second-order damping times for neutron stars, confirming that all selected equations of state (EoSs)—including nucleons-only matter, strange matter, and neutron decay into dark matter—exhibit a universal behavior. This universality is evident from the precise overlap of their plots for both first- and second-order equations. Furthermore, Fig. (3) highlights that dark matter, when incorporated into the neutron star’s composition, consistently exhibits universal behavior. This universality holds true regardless of the strength of dark matter’s self-interaction, as indicated by recent findings [44]. These results provide critical insights into neutron star dynamics across various theoretical models, offering potential constraints on the properties of dark matter and its interaction mechanisms within compact astrophysical objects. Additionally, the second-order damping time exhibits a much steeper decline compared to the first-order, with both damping times decreasing as the compactness of the neutron star increases.
The relationship between normalized damping time and neutron star mass, as depicted in Fig. (4), highlights that heavier neutron stars experience reduced damping time. Interestingly, while nucleons-only EoS and all self-interacting dark matter EoSs follow a universal trend and nearly overlap, strange matter EoS and neutron decay into dark matter EoS without self-interaction diverge from this universality.
@@ 7.2 Mass-scaled angular frequency
Figures (5) and (6) which are calculated using Ref. [122], illustrate the relationship between mass-scaled angular frequency and both compactness and effective compactness. These figures provide valuable insights into how the angular frequency varies with neutron star structure, highlighting key trends influenced by mass and gravitational binding. As shown in Fig. (5), the angular frequency of neutron stars increases with their compactness, suggesting that more compact neutron stars may emit stronger gravitational waves, potentially detectable. However, the strange matter EoS and dark matter EoS without self-interaction appear to deviate from the trend observed in nucleons-only and self-interacting dark matter EoSs. Notably, nucleons-only and self-interacting dark matter EoSs exhibit a universal behavior, aligning closely across different models.
Figure (6) illustrates the universal behavior observed across the selected equations of state (EoSs), including strange matter and neutron decay into dark matter EoS. The plot demonstrates that, despite differences in composition, these EoSs follow a consistent trend, reinforcing the idea that the mass-scaled angular frequency maintains a predictable relationship with effective compactness (). This universal pattern suggests that key physical properties governing neutron star oscillations remain largely unaffected by variations in the underlying matter composition, except for minor deviations. The alignment of different models further supports the robustness of this behavior across a wide range of astrophysical scenarios.

### v2

@@ Abstract
Furthermore, fundamental mode (f-mode) oscillations have been analyzed, which serve as a diagnostic tool for probing neutron star interiors. The f-mode frequencies follow universal relations, reinforcing their applicability for constraining dense matter properties. It has been shown that neutron stars composed of nucleons-only and self-interacting dark matter exhibit a universal behavior in damping time and angular frequency, whereas strange matter and non-self-interacting dark matter deviate from this trend.
These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer. In particular, the f-mode frequencies generated by such stars fall within the detection range of modern telescopes, making it possible to observe these signals directly. This opens a promising pathway for probing dark matter indirectly—using neutron stars as natural laboratories for fundamental physics.
@@ 1 Introduction
The hypothesis of self-interacting dark matter (SIDM) has gained prominence due to its potential to address discrepancies within the standard collisionless cold dark matter (CCDM) paradigm, alongside its theoretical plausibility within extensions of the Standard Model. Specifically, the CCDM model faces challenges such as the observed flat density profiles in dwarf galaxy cores, contrasting with the cuspy profiles predicted by simulations [1, 2], the overabundance of simulated satellite galaxies compared to observed counts [1, 3, 4], and the ”too big to fail” problem, wherein simulations predict massive, star-forming dwarf galaxies that remain undetected [5, 6].
@@
Additionally, the neutron lifetime anomaly, where beam experiments yield longer lifetimes than bottle experiments [12, 13, 14, 15], suggests a potential decay channel into undetected dark sector particles [16, 17, 18, 19]. This has motivated the exploration of neutron decay into dark matter as a potential resolution. The neutron decay scenario, proposed by Fornal and Grinstein [18], posits that neutrons decay into a dark fermion (), potentially explaining the observed neutron lifetime discrepancy. This hypothesis implies that neutron stars, composed primarily of neutrons, could serve as natural laboratories for studying dark matter interactions.
@@
In this study, the neutron decay into dark matter particle model is tested using well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear equation of state (EoS), and which have recently been revisited [28, 29]. This investigation is further extended to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
This paper has been divided into different sections. The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints. Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena. Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications. Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity. Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations. Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships. This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms. Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance. Section 8 provides a comparative analysis of f-mode frequencies across various neutron star models, offering insights into their structural and dynamical characteristics. Section 9 concludes the article by summarizing the key findings and discussing their broader relevance to neutron star physics and gravitational wave astronomy.
@@ 2 Neutron decay into dark matter
This hypothesis posits that weakly interacting dark matter particles remain undetected in beam experiments but contribute to the total decay rate in bottle experiments. This alternative to neutron-mirror neutron oscillation [16, 17] necessitates a dark fermion with a mass (mχ) nearly degenerate with the neutron within the range , more details of the model can be found in Ref. [16, 17]. The accompanying boson cannot be a photon as shown in Ref. [39, 40].
Furthermore, this dark decay hypothesis has been suggested as a potential resolution to the reactor anti-neutrino anomaly [40]. Notably, it imposes constraints on neutron star properties. If these dark fermions are non-interacting, the maximum neutron star mass is predicted to be below 0.7 M⊙ [41, 42, 43, 44, 45], contradicting observed neutron star masses exceeding 2 M⊙ [46, 47, 48, 49]. Therefore, a repulsive self-interaction among dark fermions is required to reconcile this discrepancy.
This study investigates the implications of this dark decay channel on neutron star properties. This study takes forward the previous findings and constrains [50, 17, 51, 42, 16] the properties of the dark matter using the gravitational wave, which can be very crucial in determining the interior of the neutron star.
@@ 3 Neutron Stars
The hypothesis that neutrons may decay into DM suggests that a substantial number of dark matter particles may accumulate within neutron stars, assuming dark matter is annihilating, potentially leading to observable alterations in their macroscopic properties [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 44, 97, 98, 99, 100, 101, 102, 103, 104]. To model the nuclear matter inside neutron stars with precision, the equation of state (EoS) based on the quark-meson coupling (QMC) model has been used [105, 106, 107, 108]. This model is rooted in a relativistic mean field. This framework, grounded in relativistic mean-field theory, effectively represents nucleon interactions through meson exchange. In this study, is limited to neutron decay into dark matter, hadronic composition to nucleons and strange quark matter only.
@@ 4 Equation of State
To determine the equation of state, the QMC model has been adapted, which has shown promising results in [85, 50, 109, 110]. The quark-meson coupling (QMC) model was first introduced by Guichon and later expanded upon by Guichon, Thomas, and their collaborators. In this framework, nucleons are modeled as clusters of three quarks confined within an MIT bag, emphasizing their internal structure—unlike other models that treat nucleons as point-like particles [111]. The interactions between baryons arise from meson exchange, with the mesons coupling dynamically to the confined quarks. A key feature of this model is the strong scalar mean field, which induces significant modifications in the properties of bound baryons. As proposed by Fornal and Grinstein [18, 17], if neutron stars decay into dark matter. Dark matter must be taken into account when formulating an equation of state. At ()-equilibrium, the chemical equilibrium equations governing the composition of matter are [42, 112, 76]:
@@
where is the mass of corresponding particle (subscripted). The nucleon effective mass, , given by [108, 109]
@@
The nucleons-only EoS is taken from [108] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
where the energy density of the strange matter and is the bag constant, taken to be 1014gm/fm3 [113].
@@ 6 Relationship between energy density and pressure
6.1 Effective Lagrangian for neutron–dark matter conversion
To model the possibility of neutron decay into dark matter inside neutron stars, we extend the nucleonic sector (described here using the QMC model) by including a dark fermion carrying baryon number. In order to reproduce the repulsive self-interaction assumed in the EoS analysis, we introduce a dark vector mediator , analogous to the meson in relativistic mean-field models. The effective Lagrangian reads
where . The last term represents an effective in-medium mixing between neutrons and dark fermions, which allows for neutron conversion into dark matter when . Baryon number is conserved if carries baryon number .
At scales below , integrating out the heavy vector yields a four-fermion operator,
where the effective coupling is defined as . In mean-field approximation, this produces a repulsive contribution to the energy density and pressure, , , which stiffens the equation of state as increases [41, 42, 50]. This mechanism directly explains why the dark-matter–induced EoS approaches the nucleon-only EoS for large , as shown in Figure (1), which presents the equations of state (EoS) for nucleonic matter, strange matter, and scenarios involving neutron decay into dark matter under varying dark matter self-interactions. For the nucleon-only case, the Quark-Meson Coupling (QMC) model is employed. The EoS describing neutron decay into dark matter is shown for several values of the dark matter–dark matter (DM–DM) self-repulsion strength. For comparison, the strange matter EoS is also included.
The results indicate that the nucleon-only EoS is the stiffest among the considered models. As the DM–DM self-repulsion increases, the corresponding EoS for neutron decay into dark matter becomes progressively stiffer and approaches the nucleon-only case which is in agreement with the findings of [50, 41, 51]. This behavior arises because stronger DM–DM self-interaction raises the energy cost for neutron decay into dark matter, thereby reducing the overall dark matter population within the system [50]. Consequently, larger self-interaction strengths, , result in fewer dark matter particles and stiffer equation of state [50].
The strange matter EoS provides a poor description of nuclear matter at low energy densities. As shown in Fig. (1), it coincides with the nucleon-only EoS up to approximately 500 MeV/fm3 [108, 109, 50]. Beyond this density, strange matter begins to manifest, and the corresponding EoS deviates from both the nucleon-only and the neutron-decay EoSs. In this regime, the strange matter EoS is comparatively softer than the other two models. The deviation of the strange matter EoS at densities above MeV/fm3 arises from the bag constant assumption in the MIT bag model, which lowers the pressure support compared to nucleonic models. As a result, the sound speed is reduced, leading to longer damping times and lower f-mode frequencies. This implies that strange stars may emit weaker gravitational waves, with potentially distinguishable signals compared to nucleonic stars of the same mass.
@@ 6.1 Effective Lagrangian for neutron–dark matter conversion
Fig. (2) represents the mass and radius of neutron stars based on different EoSs. Ref. [119, 120, 121] report the observation of neutron stars of a mass of exceeding 2 M⊙ [46, 47, 48, 49], indicating that any viable model must support a maximum neutron star mass above this threshold 2 M⊙. The nucleon-only EoS produces a neutron star of maximum mass greater than 2 M⊙, the strange matter EoS produces a neutron star of maximum mass close to 2 M⊙, while different EoSs based on different DM-DM self-interactions produce a neutron star of less than 2 M⊙. The non-self-interactive ( = 0 fm2) DM EoS produces a neutron star of mass close to 0.7 M⊙ which is well below the acceptable value and alignes with [41, 17, 51]. When the DM-DM self-interaction [50] the neutron decay into dark matter EoS produces a neutron star of 2 M⊙. Therefore, acceptable self-interaction strength is 26 fm2. The failure of the fm2 case to support neutron stars above highlights that non-self-interacting dark fermions dilute pressure support too severely. Physically, the DM particles form a diffuse distribution, decreasing compactness and oscillation frequencies.
@@ 7 Universalities
The universalities refer to empirical relations that link the fundamental oscillation modes of compact objects, such as neutron stars and dark stars, to their macroscopic properties like mass, radius, and compactness. These universal relations are valuable because they provide a model-independent way to infer stellar characteristics from gravitational wave observations [122, 117]. Traditional neutron stars follow well-established scaling laws for f-mode frequencies and damping times, but recent studies show that dark matter admixed compact stars deviate from these relations, requiring new frameworks to describe their oscillations [123, 124]. By leveraging f-mode gravitational wave signals, astrophysicists aim to refine equations of state and probe exotic forms of matter beyond the Standard Model. The detailed analysis on tidal deformability of neutron decay mode is given in [50].
@@ 7.1 Damping time
The damping time of f-mode oscillations is related to the balance between the total mode energy and the rate of energy loss via gravitational-wave emission ,
The oscillation energy for a mode with angular frequency is given by
where is the Lagrangian displacement vector and the local energy density. The gravitational-wave luminosity for the dominant quadrupole mode is
From these relations, the universal scaling emerges [125], which we adopt and extend in this study.
Reference [28] introduced a first-order universal relation for the damping time () of f-mode oscillations, derived from the quadrupole radiation formula. The relation, shown in Figure (3), is expressed as
Here reflects the ratio of oscillation energy to gravitational-wave luminosity, implying that more massive stars radiate more efficiently and exhibit shorter damping times. Lioutas and Stergioulas [126] later extended this relation with higher-order corrections, confirming the monotonic decrease of with mass. In comparison, strange matter stars exhibit longer damping times than both nucleon-only stars and neutron stars with dark matter cores.
@@ 7.2 Mass-scaled angular frequency
Figures (4) and (5) illustrate the relation between the mass-scaled angular frequency, where , and two structural parameters of neutron stars: compactness () and [117]. These diagnostics are particularly valuable since they encode information about the distribution of mass and the role of internal composition in determining stellar oscillation properties. The parameter , in particular, is closely tied to the efficiency of gravitational-wave (GW) emission, as it relates the global stellar mass to the moment of inertia, thereby providing insight into how matter is distributed inside the star.
As Figure (4) shows, when 26 fm2, the nucleon-only and self-interacting DM EoSs follow a universal scaling curve, independent of the specific interaction strength. This suggests that compactness is a reliable predictor of oscillation properties, robust against moderate variations in microphysics. Strange matter and non-self-interacting DM deviate significantly. The strange matter deviation reflects the reduced stiffness of the EoS, while the DM case is destabilized by the large DM population at the core, altering the star’s effective compactness. These deviations may serve as diagnostic signatures of exotic compositions.
Figure (5) further reinforces the presence of a quasi-universal relation across different microscopic models when the angular frequency is expressed in terms of . Remarkably, even when exotic compositions are considered—including strange matter and neutron-decay into DM with strong self-interaction ( fm2)—the data points collapse onto a nearly single curve. This indicates that effectively captures the essential structural scaling of neutron stars, smoothing out differences in composition. Deviations are observed only in the dark matter without self-repulsion case, but the overall correlation persists. This reinforces the robustness of universal relations in gravitational-wave asteroseismology.